# A Genealogy of Foundation Models in Remote Sensing

### v2

@@ 1. Introduction
With the recent successes of Natural Language Processing (NLP) frameworks like Generative Pre-trained Transformers (GPT) (OpenAI et al., 2024) and Bi-directional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), the development and application of machine learning is shifting away from focusing on transfer learning. Transfer learning uses traditional supervised learning on a large, labeled dataset to learn representations of that data to apply to other related tasks. Instead, Self-Supervised Learning (SSL), which learns general-purpose representations of data without the need for external labels, has risen in popularity (Chen et al., 2024). As datasets have grown larger, with some datasets now containing billions of data points (Dehghani et al., 2023), the process of labeling data for supervised learning has proven to be a costly bottleneck for improving model performance. SSL circumvents this issue with the use of pretext tasks: auxiliary learning objectives designed to leverage the natural structure of unlabeled data in order to learn patterns or relationships within the data. Researchers can then fine-tune these pre-trained models for specific downstream tasks on significantly smaller labeled datasets. Fine-tuning enables downstream users to achieve state-of-the-art (SOTA) performance on a specific task while incurring significantly lower costs than training an equivalent model from scratch (Anisuzzaman et al., 2025). Thus, effective SSL training democratizes access to high performance machine learning models.
@@ 2. Prior Work
Recent papers in the field have surveyed available foundation models in RS. Zhou explored the high potential of applied Vision Language Models (VLMS) while highlighting their few shot capabilities in downstream tasks (Zhou et al., 2024). Chen inspected how models can adapt to utilize non-raster based data (Chen et al., 2024), such as poly-lines, to represent transportation networks. Zhang examined the singular task of building foundation models specifically for urban environments (Zhang et al., 2024). Mai created a ’wish list’ of desirable traits in RS foundation models, with rationale for each wish, and discussed where current foundation models exceed or fall short of these goals (Mai et al., 2023a). Lu categorized the model backbones, training datasets, image resolution, and over-arching pre-training methods for existing RS foundation models (Lu et al., 2025). Lu also documented the number of parameters for each model and compiled results across a wide range of diverse downstream tasks (Lu et al., 2025).
@@
Figure 1. Satellite imagery of Boston, MA in October, 2020. SAR imagery is obtained from Sentinel-1, and optical imagery is obtained from Sentinel-2. SAR false color uses VV polarization for red channel, VH polarization for green channel, and the average of VV and VH polarizations for blue channel. NDVI refers to the Normalized Difference Vegetation Index, an index based on the Red and Near Infrared (NIR) bands that is utilized to assess vegetation health.:
@@ 4.1. Negative Sampling Models
Figure 2. CV foundation model architectures that leverage contrastive learning.:
@@ 4.2. Distillation Frameworks
Distillation simply refers to the concept of one model being trained on the outputs of another model (Phuong and Lampert, 2019). Distillation is not strictly an SSL method, as some research focuses on using distillation to ’compress’ knowledge from a large pre-trained model into a smaller model (Wu et al., 2024). Distillation leverages two encoders, dubbed the “teacher” and the “student”. When trying to ’compress’ logic, the teacher is some pre-trained model, but in the context of SSL, both student and teacher encoders are learning, albeit not at the same rate or with the same method, depending on the specific set up. Both are fed in augmented versions of an image to output embeddings. Then, contrastive loss is performed between the teacher and student embeddings and back propagated through the student network. The teacher network typically learns via the weighted average update system that was introduced in MoCo (He et al., 2020). Distillation models such as Joint-Embeddings Predictive Architecture (JEPA) (Assran et al., 2023) build on this framework by introducing a decoder after the student to predict the embeddings from the teacher. This approach leverages reconstruction loss in the latent space, but more distillation SSL approaches rely on contrasting embeddings from the student and teacher encoders (Grill et al., 2020; Chen et al., 2022; Jang et al., 2023; Oquab et al., 2024).
@@ 4.3. Redundancy Reduction
Figure 3. An illustration of a CV foundation model that relies on redundancy reduction.:
@@ 4.4. Masked Image Modeling
Figure 4. CV foundation model framework for masked image modeling:
@@ 5.2. Geolocation Pairing
SatCLIP (Klemmer et al., 2024) also follows the CLIP approach while utilizing spatial encodings as one of its modalities and optical imagery captured from Sentinel-2 as the other. SatCLIP takes in all Sentinel-2 bands and does not limit itself purely to RGB data. SatCLIP does not utilize Random Fourier Features for the spatial encodings, and instead uses a SirenNet (Sitzmann et al., 2020) that relies on spherical harmonics (Rußwurm et al., 2024) to learn effective spatial encodings.
Geo-Aligned Implicit Representations (GAIR) (Liu et al., 2025) takes this a step further by utilizing location as a way to align multispectral satellite images from Sentinel-2 and ground-level RGB images from the Global Streetscapes dataset (Hou et al., 2024). Ground-level imagery can provide a highly informative mode to train satellite imagery against. However, a single satellite image’s scope is so large that hundreds or thousands of ground-level images can correspond to a single satellite image. In order to address this issue from a contrastive learning standpoint, GAIR learns Implicit Neural Representations (INR) (Xu et al., 2022) of a Sentinel-2 image. INR learns a continuous function for spatial coordinates to corresponding signals. These INR encodings of Sentinel-2 imagery are then contrasted against location encodings and ground level vision encodings. GAIR uses RFF (Tancik et al., 2020) for location encodings and a standard ViT for ground-level image encodings.
@@ 5.4. Image-Text Pairing
Figure 5. An illustration of shared and unique concepts among RS foundation models which utilize contrastive learning with negative sampling for their SSL task.:
@@ 6.2. Multiple Contrastive Tasks
After these contrastive tasks, the embeddings go through a multi-modal temporal fusion encoder at both the student and teacher level. This results in a fused embedding. Sinkhorn-Knopp is applied again, but this time, the resulting clusters are compared to geographic regions in order to learn geo-context. Then the fused embeddings go through the same contrastive learning tasks described in the paragraph above. The resulting model was reported to achieve SOTA performance on downstream tasks.
Figure 6. An illustration of shared and unique concepts among RS foundation models which utilize knowledge distillation for their SSL task.:
@@ 7. Redundancy Reduction in Remote Sensing
DeCUR’s key distinction from other multi-modal approaches we have discussed lies in its desire to learn inter-modal representations, but preserve data unique to a given modality. It does so by setting aside a percentage of the latent space for shared representations and leaving the other percentage of the latent space for modality specific information. When the Barlow Twin’s redundancy reduction loss is applied from one mode to another, DeCUR only calculates this loss for shared dimensions of the latent space. DeCUR also calculates redundancy reduction loss within a given mode itself, and does so for all of the dimensions within the latent space. This strategy enables DeCUR to align modalities without losing modality-specific information and results in SOTA performance on downstream tasks.
@@ 8.1. Masking Approach
Spatial-Spectral MAE (S2MAE) (Li et al., 2024a) takes this a step further, as they find that masking 90% of satellite imagery yields the best results in their novel framework. S2MAE trains on multispectral imagery from FMoW (Christie et al., 2018) and BigEarthNet (Sumbul et al., 2019) datasets. Unlike RingMo and other image RGB masking strategies which mask out the same pixel across all channels, S2MAE employs independent 3D masking, which masks out different pixels in each channel. 3D masking was introduced for videos and appears to learn environmental details well, but struggles to track movement from frame to frame (Feichtenhofer et al., 2022). While this is poorly suited for video analysis, it is ideal for RS applications, as seen in S2MAE’s SOTA performance on the Onera Satellite Change Detection (OSCD) (Daudt et al., 2018) and EuroSat (Helber et al., 2019) benchmarks. S2MAE also finds that shallow decoders are poorly suited for MAE with multispectral imagery, a notable departure from established CV practices (He et al., 2021).
@@ 8.2. Temporal Encodings
Presto (Tseng et al., 2024) takes this approach even further, constructing a pixel-timeseries training dataset that consists of input from multiple types of data, including Sentinel-2 RGB, Sentinel-1 SAR, DEM, and the Dynamic World (Brown et al., 2022) dataset. Dynamic World contains land cover labels for 10m resolution Sentinel-2 imagery. The framework then employs a variety of different strategies for training. It masks channels, blocks of timesteps, single timesteps, and/or random pixels for reconstruction by a downstream encoder. The resulting encoder performs well on feature extraction, fine-tuning, and time series forecasting. The variety of masking strategies ensures that Presto can successfully process time series datasets with missing channels, a variety of temporal resolutions, and only a small subset of timesteps. However, performance on tasks that are temporally agnostic and rely on a single image lags behind SOTA models. The authors note that this is possibly due to the model’s inability to process images with a spatial resolution higher than 10m.
@@ 8.3. Spatial Resolution
Where ScaleMAE focuses on understanding spatial resolution of the overall image, USat (Irvin et al., 2023) chooses to focus on spatial resolution across image channels. Different bands of the same sensor can have dramatically different spatial resolutions. For instance, in the case of Sentinel-2, bands (i.e., input channels) range from 10m resolution to 60m resolution. Many approaches up- or down- sample channels to yield a consistent spatial resolution, and some simply drop low resolution channels altogether (Jain et al., 2022a). This ensures that patch sizes remain consistent, but may lose resolution specific information. USat trains on Sentinel-2 and NAIP, and adjusts the number of patches per channel depending on the spatial resolution. High resolution channels have more patches to capture specific high resolution information. Lower resolution channels have fewer patches, but the total number of patches must be evenly divisible by the number of high resolution patches. This enables channels with different resolutions to cleanly map to one another. Each channel will be masked by the same percentage. More patches are masked in high resolution channels to preserve the percentage of the image being masked. Then, patches are aggregated via spectral pooling, and a standard decoder is used to reconstruct the original image. This approach effectively extracts high-level and low-level features from differing resolutions, but does have some drawbacks. Namely, the assumption that different sensors will have a resolution that is cleanly divisible will not always be true. The authors also note that the approach becomes costly to train with multiple different sensors.
@@
Figure 7. An illustration of shared and unique concepts among RS foundation models which utilize masked auto-encoding for their SSL task and train on multispectral imagery from Sentinel-2 and potentially some other multi-sensor data.:
@@ 8.5. Orientation Awareness
Figure 8. An illustration of shared and unique concepts among RS foundation models which utilize masked image modeling for their SSL task and all train only on RGB images.:
@@ 8.6. Wavelength Awareness
Dynamic One-For-All (DOFA) (Xiong et al., 2024) attempts to address this problem by creating a framework that can take in data from any number of different sensors. DOFA follows a hypernetwork approach (Chauhan et al., 2024), where a hypernetwork takes in the wavelength details of a given band and then generates weights that are used to initialize a neural network. Image bands are masked, then fed into that neural network to generate encodings. The encodings are then fed through an encoder-decoder structure, with a mirrored hypernetwork at the end responsible for reconstructing the original image. This framework is designed to learn bandwidth specific features via the hypernetwork and shared features via the shared encoder-decoder. DOFA is trained on SAR imagery from Sentinel-1, multispectral imagery from Sentinel-2, hyperspectral imagery from enMAP, and high resolution RGB imagery from NAIP. The resulting framework achieves SOTA performance on downstream tasks such as BigEarthNet (Sumbul et al., 2019), So2Sat (Zhu et al., 2019), and EuroSat (Helber et al., 2019).
Foundation Models for Forest Monitoring (FoMo) (Bountos et al., 2025) also trains to be aware of spectral wavelength, but does so without the use of a hypernetwork. FoMo trains on four satellite-based datasets and four aerial-based dataset, with spatial resolutions ranging from less than 5cm to 60m. The datasets include MS imagery from Landsat 8-9, Sentinel-2, and Gaofen-2. They also include SAR imagery from Sentinel-1, DEM data, and both RGB and MS imagery from high-res UAV sensors. The resulting training dataset is rich in spatial and spectral diversity. FoMo’s flexible training strategy groups together all of the images of the same spatiotemporal scene, then randomly selects bands to use in that particular sample. Each band is independently tokenized to avoid spatial resolution issues. FoMo then incorporates standard positional embeddings and introduces spectral embeddings to denote what band each channel belongs to. This setup allows for learned representations across bands and within bands. Tokens are masked, passed through a ViT encoder, and then go through a standard decoder for reconstruction. Due to the random band selection process for generating samples, batches can potentially be very different. Gradient accumulation across batches is used to ’smooth’ the learning rate.

### v3

@@ 1. Introduction
With the recent successes of Natural Language Processing (NLP) frameworks like Generative Pre-trained Transformers (GPT) (OpenAI et al., 2024) and Bi-directional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), the development and application of machine learning is shifting away from traditional supervised learning. Supervised learning on a large, labeled dataset learns representations of that data. These representations are most applicable to the task they have trained on, but can be transferred to other similar tasks to speed up training. Instead, Self-Supervised Learning (SSL), which learns general-purpose representations of data without the need for external labels, has risen in popularity (Chen et al., 2024). As datasets have grown larger, with some datasets now containing billions of data points (Dehghani et al., 2023), the process of labeling data for supervised learning has proven to be a costly bottleneck for improving model performance. SSL circumvents this issue with the use of pretext tasks: auxiliary learning objectives designed to leverage the natural structure of unlabeled data in order to learn patterns or relationships within the data. Researchers can then fine-tune these pre-trained models for specific downstream tasks on significantly smaller labeled datasets. Fine-tuning enables downstream users to achieve state-of-the-art (SOTA) performance on a specific task while incurring significantly lower costs than training an equivalent model from scratch (Anisuzzaman et al., 2025). Thus, effective SSL training democratizes access to high performance machine learning models.
@@ 2. Prior Work
Recent papers in the field have surveyed available foundation models in RS. Zhou et al. explored the high potential of applied Vision Language Models (VLMS) while highlighting their few shot capabilities in downstream tasks (Zhou et al., 2024). Chen et al. inspected how models can adapt to utilize non-raster based data (Chen et al., 2024), such as poly-lines, to represent transportation networks. Zhang et al. examined the singular task of building foundation models specifically for urban environments (Zhang et al., 2024). Mai et al. created a ’wish list’ of desirable traits in RS foundation models, with rationale for each wish, and discussed where current foundation models exceed or fall short of these goals (Mai et al., 2023a). Lu et al. categorized the model backbones, training datasets, image resolution, and over-arching pre-training methods for existing RS foundation models (Lu et al., 2025). Lu et al. also documented the number of parameters for each model and compiled results across a wide range of diverse downstream tasks (Lu et al., 2025).
@@
Figure 1. Satellite imagery of Boston, MA in October, 2020. SAR imagery is obtained from Sentinel-1, and optical imagery is obtained from Sentinel-2. SAR false color uses VV polarization for red channel, VH polarization for green channel, and the average of VV and VH polarizations for blue channel. NDVI refers to the Normalized Difference Vegetation Index, an index based on the Red and Near Infrared (NIR) bands that is utilized to assess vegetation health.
A variety of satellite images capturing Boston, MA in the month of October, 2020. Images with VV and VH polarizations captured by Synthetic Aperture Radar (SAR) are in black and white and do a good job of capturing water and urban structures. A false color image of SAR combines the different polarizations to provide a sharper, 3-channel image. There is a standard RGB optical image, and a false color image with Shortwave Infrared (SWIR) bands and the blue light band that is used to highlight geological features. Then, there is a depiction of NDVI, where dark green highlights healthy vegetation, dark blue highlights no vegetation, and white indicates sparse or unhealthy vegetation.
@@ 3.1. Optical Imagery
Many satellite missions employ mission-specific data pre-processing pipelines and provide multiple standardized data products. For example, Sentinel-2 offers two primary optical products, and different teams have chosen either to train foundation models:
  • Level-1C (L1C) provides Top-of-Atmosphere (TOA) reflectance for all 13 spectral bands and includes standard radiometric calibration, geometric correction, and ortho-rectification, but does not apply atmospheric correction.(Spoto et al., 2012)
  • Level-2A (L2A) provides Bottom-of-Atmosphere (BOA), or surface reflectance, imagery derived via atmospheric correction. This product excludes B10, which is not suitable for surface reflectance retrieval.(Main-Knorn et al., 2017)
The majority of remote-sensing foundation models discussed in this manuscript that leverage Sentinel-2 data are trained using the L1C product. Unless otherwise specified, models trained on Sentinel-2 data are assumed to use L1C imagery.
@@ 4.1. Negative Sampling Models
Figure 2. CV foundation model architectures that leverage contrastive learning.
Two small diagrams that depict contrastive learning with negative sampling or distillation. In negative sampling, we see a batch of images augmented, passed through the same encoder, and then compared. In distillation, we see a single image augmented, passed through two different encoders, and then compared.
@@ 4.2. Distillation Frameworks
Distillation simply refers to the concept of one model being trained on the outputs of another model (Phuong and Lampert, 2019). Distillation is not strictly an SSL method, as some research focuses on using distillation to ’compress’ knowledge from a large pre-trained model into a smaller model (Wu et al., 2024). Distillation leverages two encoders, dubbed the “teacher” and the “student”. When trying to ’compress’ logic, the teacher is some pre-trained model, but in the context of SSL, both student and teacher encoders are learning, albeit not at the same rate or with the same method, depending on the specific set up. Both are fed in augmented versions of an image to output embeddings. Then, contrastive loss is performed between the teacher and student embeddings and back propagated through the student network. The teacher network typically learns via the weighted average update system that was introduced in MoCo (He et al., 2020). The above broadly describes the framework of BYOL (Grill et al., 2020). DINO (Caron et al., 2021) has a very similar framework, but utilizes basic centering and sharpening operations on the output from the teacher to avoid embedding collapse. Distillation models such as Joint-Embeddings Predictive Architecture (JEPA) (Assran et al., 2023) build on this framework by introducing a decoder after the student to predict the embeddings from the teacher. This approach leverages reconstruction loss in the latent space, but more distillation SSL approaches rely on contrasting embeddings from the student and teacher encoders (Grill et al., 2020; Chen et al., 2022; Jang et al., 2023; Oquab et al., 2024).
@@ 4.3. Redundancy Reduction
Figure 3. An illustration of a CV foundation model that relies on redundancy reduction.
A diagram that depicts learning via redundancy reduction. A single image is augmented, and passed through an encoder to generate two encodings. A cross-correlation matrix is calculated, and then compared to the identity matrix for learning.
@@ 4.4. Masked Image Modeling
Figure 4. CV foundation model framework for masked image modeling
A diagram depicting masked image modeling, where a given image is obscured by a large amount, and then passed through an encoder to generate an embedding. That embedding is then passed through a decoder which attempts to recreate the original image.
@@ 5.2. Geolocation Pairing
SatCLIP (Klemmer et al., 2024) also follows the CLIP approach while utilizing spatial encodings as one of its modalities and optical imagery captured from Sentinel-2 as the other. SatCLIP does not limit itself purely to RGB data, and rather trains on the 12 bands of data present in Sentinel-2’s L2A data product. SatCLIP does not utilize Random Fourier Features for the spatial encodings, and instead uses a SirenNet (Sitzmann et al., 2020) that relies on spherical harmonics (Rußwurm et al., 2024) to learn effective spatial encodings.
Geo-Aligned Implicit Representations (GAIR) (Liu et al., 2025) takes this a step further by utilizing location as a way to align multispectral satellite images from Sentinel-2 and ground-level RGB images from the Global Streetscapes dataset (Hou et al., 2024). As with SatCLIP, GAIR trains on Sentinel-2’s L2A data product. Ground-level imagery can provide a highly informative mode to train satellite imagery against. However, a single satellite image’s scope is so large that hundreds or thousands of ground-level images can correspond to a single satellite image. In order to address this issue from a contrastive learning standpoint, GAIR learns Implicit Neural Representations (INR) (Xu et al., 2022) of a Sentinel-2 image. INR learns a continuous function for spatial coordinates to corresponding signals. These INR encodings of Sentinel-2 imagery are then contrasted against location encodings and ground level vision encodings. GAIR uses RFF (Tancik et al., 2020) for location encodings and a standard ViT for ground-level image encodings.
@@ 5.4. Image-Text Pairing
Figure 5. An illustration of shared and unique concepts among RS foundation models which utilize contrastive learning with negative sampling for their SSL task.
A sort of Venn diagram where a variety of RS models that utilize contrastive learning with negative sampling are compared with one another to highlight similarities and differences.
@@ 6.2. Multiple Contrastive Tasks
After these contrastive tasks, the embeddings go through a multi-modal temporal fusion encoder at both the student and teacher level. This results in a fused embedding. Sinkhorn-Knopp is applied again, but this time, the resulting clusters are compared to geographic regions in order to learn geo-context. Then the fused embeddings go through the same contrastive learning tasks described in the paragraph above. The resulting model was reported to achieve SOTA performance on a plethora of downstream tasks at the time of publication in 2024. These tasks included:
  • Segmentation tasks, such as Potsdam (Sherrah, 2016)
  • Object detection tasks, such as DIOR(Li et al., 2020)
  • Change detection tasks, such as OSCD(Daudt et al., 2018)
  • Classification tasks, such as MillionAID(Long et al., 2021)
Figure 6. An illustration of shared and unique concepts among RS foundation models which utilize knowledge distillation for their SSL task.
A sort of Venn diagram where a variety of RS models that utilize distillation SSL are compared with one another to highlight similarities and differences.
@@ 7. Redundancy Reduction in Remote Sensing
DeCUR’s key distinction from other multi-modal approaches we have discussed lies in its desire to learn inter-modal representations, but preserve data unique to a given modality. It does so by setting aside a percentage of the latent space for shared representations and leaving the other percentage of the latent space for modality specific information. When the Barlow Twin’s redundancy reduction loss is applied from one mode to another, DeCUR only calculates this loss for shared dimensions of the latent space. DeCUR also calculates redundancy reduction loss within a given mode itself, and does so for all of the dimensions within the latent space. This strategy enables DeCUR to align modalities without losing modality-specific information and obtained SOTA performance on BigEarthNet-MM (Sumbul et al., 2019) classification tasks at the time of publication in 2024.
@@ 8.1. Masking Approach
Spatial-Spectral MAE (S2MAE) (Li et al., 2024a) takes this a step further, as they find that masking 90% of satellite imagery yields the best results in their novel framework. S2MAE trains on multispectral imagery from FMoW (Christie et al., 2018) and BigEarthNet (Sumbul et al., 2019) datasets. Unlike RingMo and other image RGB masking strategies which mask out the same pixel across all channels, S2MAE employs independent 3D masking, which masks out different pixels in each channel. 3D masking was introduced for videos and appears to learn environmental details well, but struggles to track movement from frame to frame (Feichtenhofer et al., 2022). While this is poorly suited for video analysis, it is ideal for RS applications, as seen in S2MAE’s SOTA performance on the Onera Satellite Change Detection (OSCD) (Daudt et al., 2018) and EuroSat (Helber et al., 2019) benchmarks at the time of publication in 2024. S2MAE also finds that shallow decoders are poorly suited for MAE with multispectral imagery, a notable departure from established CV practices (He et al., 2021).
@@ 8.2. Temporal Encodings
Presto (Tseng et al., 2024) takes this approach even further, constructing a pixel-timeseries training dataset that consists of input from multiple types of data, including Sentinel-2 RGB, Sentinel-1 SAR, DEM, and the Dynamic World (Brown et al., 2022) dataset. Where other foundation models typically operate on patches or convolutions of pixels, Presto operates explicitly on pixels. This choice is more costly from a computational standpoint, but theoretically results in high-resolution representations. Dynamic World contains land cover labels for 10m resolution Sentinel-2 imagery. The framework then employs a variety of different strategies for training. It masks channels, blocks of timesteps, single timesteps, and/or random pixels for reconstruction by a downstream encoder. The resulting encoder performs well on feature extraction, fine-tuning, and time series forecasting. The variety of masking strategies ensures that Presto can successfully process time series datasets with missing channels, a variety of temporal resolutions, and only a small subset of timesteps. However, performance on tasks that are temporally agnostic and rely on a single image lags behind SOTA models. The authors note that this is possibly due to the model’s inability to process images with a spatial resolution higher than 10m.
@@ 8.3. Spatial Resolution
Where ScaleMAE focuses on understanding spatial resolution of the overall image, USat (Irvin et al., 2023) chooses to focus on spatial resolution across image channels. Different bands of the same sensor can have dramatically different spatial resolutions. For instance, in the case of Sentinel-2, bands (i.e., input channels) range from 10m resolution to 60m resolution. Many approaches up- or down- sample channels to yield a consistent spatial resolution, and some simply drop low resolution channels altogether (Jain et al., 2022a). This ensures that patch sizes remain consistent, but may lose resolution specific information. USat trains on Sentinel-2 and NAIP, and adjusts the number of patches per channel depending on the spatial resolution. High resolution channels have more patches to capture specific high resolution information. Lower resolution channels have fewer patches, but the total number of patches must be evenly divisible by the number of high resolution patches. This enables channels with different resolutions to cleanly map to one another. Each channel will be masked by the same percentage. More patches are masked in high resolution channels to preserve the percentage of the image being masked. Then, patches are aggregated via spectral pooling, and a standard decoder is used to reconstruct the original image. This approach effectively extracts high-level and low-level features from differing resolutions, but does have some drawbacks. Namely, the assumption that different sensors will have resolutions that are evenly divisible will not always be true. The authors also note that the approach becomes costly to train with multiple different sensors.
@@
Figure 7. An illustration of shared and unique concepts among RS foundation models which utilize masked auto-encoding for their SSL task. These models all encode multispectral imagery from Sentinel-2 and potentially some other multi-sensor data.
A sort of Venn diagram where a variety of remote sensing masked auto-encoding foundation models are compared and contrasted to highlight their similarities and differences. The models in question train on data acquired from multiple sensors or contain imagery captured by a multispectral sensor.
@@ 8.5. Orientation Awareness
Figure 8. An illustration of shared and unique concepts among RS foundation models which utilize masked image modeling for their SSL task and all train only on RGB images.
A sort of Venn diagram where a variety of remote sensing masked image modeling foundation models are compared and contrasted to highlight their similarities and differences. The models in question all train only on RGB imagery.
@@ 8.6. Wavelength Awareness
Foundation Models for Forest Monitoring (FoMo) (Bountos et al., 2025) trains on four satellite-based datasets and four aerial-based dataset, with spatial resolutions ranging from less than 5cm to 60m. The datasets include MS imagery from Landsat 8-9, Sentinel-2, and Gaofen-2. They also include SAR imagery from Sentinel-1, DEM data, and both RGB and MS imagery from high-res UAV sensors. The resulting training dataset is rich in spatial and spectral diversity. FoMo’s flexible training strategy groups together all of the images of the same spatiotemporal scene, then randomly selects bands to use in that particular sample. Each band is independently tokenized to avoid spatial resolution issues. FoMo then incorporates standard positional embeddings and introduces spectral embeddings to denote what band each channel belongs to. This setup allows for learned representations across bands and within bands. Tokens are masked, passed through a ViT encoder, and then go through a standard decoder for reconstruction. Due to the random band selection process for generating samples, batches can potentially be very different. Gradient accumulation across batches is used to ’smooth’ the learning rate.